OpenStudy (anonymous):

Find tan2A when cosA = -3/5 and π/2

OpenStudy (moonlitfate):

Are you sure that it's tan2A and not tanA?

OpenStudy (anonymous):

Yes I'm sure it's tan2A.

OpenStudy (moonlitfate):

Okay. Well the \[\tan \theta = \frac{ \sin \theta }{ \cos \theta }\]

OpenStudy (moonlitfate):

Since you know the cos of A is \[\frac{ -3 }{ 5}\] the sine is of A is the reciprocal of cosine: \[\frac{ -5 }{ 3 }\]

OpenStudy (btaylor):

the double angle identity for tan is: \[\large \tan(2x) = \frac{2 \tan (x)}{1 - \tan^2 (x)}\] Since we know that cos(x) is -3/5 and sin(x) is 5/3 (since it is in the second quadrant, sine is positive), we can replace tan(x) with sin(x) and cos(x). Can you do it from here?

OpenStudy (anonymous):

sin(x) =5/3 !!?!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (campbell_st):

you can't have Sin A = -5/3 ... its impossible... since the range of sin(x) is -1<= x<= 1

OpenStudy (btaylor):

sin A would be 4/5

OpenStudy (campbell_st):

draw a triangle  find x by pythagoras then you'll know sin A please... sin A -5/3 wow

OpenStudy (btaylor):

yeah, my bad...